OpenStudy (anonymous):

help...Choose the equation of the line passing through the point (-4, -2) and parallel to y = 1 fourthx + 2.

OpenStudy (anonymous):

answer choices y = 4x -1 y = 4x - 3 y = 1/4x - 3 y = 1/4x - 1

OpenStudy (anonymous):

DO you know what the slope of the line should be?

OpenStudy (anonymous):

No